2. Let P (x) = 2x4 +15x3 +31x2 +20x +4

(a) Determine whether (x −1) is a factor of P (x). (2)

(b) Find all the possible rational zeros of P (x) by using the Rational Zeros Theorem. (2)

(c) Solve P (x) = 0


1
Expert's answer
2022-05-11T12:43:12-0400

2.

(a)

"P(1)=2(1)^4 +15(1)^3 +31(1)^2 +20(1)+4"

"=72\not=0"

Therefore "(x \u22121)" is not a factor of "P (x)."

(b) Since all coefficients are integers, we can apply the Rational Zeros Theorem.

The coefficient of the constant term is 4. Find its factors: "\pm1, \pm2, \pm4."

These are the possible values for "p."

The leading coefficient is "2." Find its factors: "\pm1, \pm2."

These are the possible values for "q."

Find all possible values of "p\/q:" "\pm\dfrac{1}{2}, \pm1, \pm2, \pm4."

Possible rational roots: "\pm\dfrac{1}{2}, \pm1, \pm2, \pm4."


(c)

Check the possible roots


"\dfrac{1}{2}: P(\dfrac{1}{2})=2(\dfrac{1}{2})^4 +15(\dfrac{1}{2})^3 +31(\dfrac{1}{2})^2 +20(\dfrac{1}{2})+4"

"=\dfrac{95}{4}\not=0"


"-\dfrac{1}{2}: P(-\dfrac{1}{2})=2(-\dfrac{1}{2})^4 +15(-\dfrac{1}{2})^3 +31(-\dfrac{1}{2})^2"

"+20(-\dfrac{1}{2})+4=0"


"1: P(1)=2(1)^4 +15(1)^3 +31(1)^2 +20(1)+4"

"=72\not=0"


"-1: P(-1)=2(-1)^4 +15(-1)^3 +31(-1)^2"

"+20(-1)+4=2\not=0"


"2: P(2)=2(2)^4 +15(2)^3 +31(2)^2 +20(2)+4"

"=320\not=0"


"-2: P(-2)=2(-2)^4 +15(-2)^3 +31(-2)^2"

"+20(-2)+4=0"


"4: P(4)=2(4)^4 +15(4)^3 +31(4)^2 +20(4)+4"

"=2052\not=0"


"-4: P(-4)=2(-4)^4 +15(-4)^3 +31(-4)^2"

"+20(-4)+4=-28\not=0"

Actual rational roots: "-\dfrac{1}{2}, -2."


"2x^4+15x^3 +31x^2 +20x +4"

"=(2x+1)(x+2)(x^2+5x+2)"

"x^2+5x+2=0"

"D=25-8=17"

"x=\dfrac{-5\pm\sqrt{17}}{2}"

Total roots:"-\dfrac{1}{2}, -2,\dfrac{-5-\sqrt{17}}{2},\dfrac{-5+\sqrt{17}}{2}"